OpenStudy (anonymous):

Suppose the circumference of a circle is 2π What is its radius?

OpenStudy (igreen):

Formula: \(\sf C = 2 \pi r\) Plug in what we know: \(\sf 2 \pi = 2 \pi r\) Divide 2 pi to both sides, what's 2 pi / 2 pi?

OpenStudy (anonymous):

1 pi ?? @iGreen

OpenStudy (igreen):

Yes, but we no longer have pi.

OpenStudy (igreen):

pi / pi cancels out.l

OpenStudy (igreen):

So we have: \(\sf r = 1\)

OpenStudy (anonymous):

so its just 1 @iGreen

OpenStudy (igreen):

Yep!

OpenStudy (anonymous):

THX! @iGreen

OpenStudy (igreen):

Np!